# Understanding-in-Generation:Reinforcing Generative Capability of Unified Model via Infusing Understanding into Generation

### v2

@@ 1 Introduction
First, for verification-based reasoning (e.g., ImageCoT Zhang et al. (2025a)), as illustrated in Figure 1 (a), this method constructs multiple generative branches through repeated sampling and assesses the generative potential throughout the generation process. Specifically, the state of each branch is evaluated to determine whether it still has the potential to continue generating. If the assessment gives a negative result, the branch is terminated. Here, the understanding of the unified model is applied exclusively to the verification of intermediate images, selecting the best output from among numerous samples. However, in this framework, the understanding capability is employed merely as a tool for validation and filtering, rather than providing effective guidance during the generation. As a result, the generative ability remains confined to scopes that can be reached through repeated sampling. As shown in Figure 1 (a), given the input prompt, [”There is a cup positioned behind the woman”], all outputs consistently place the cup in front of the woman. Consequently, even the best result obtained through verification-based reasoning fails to satisfy the spatial relationship specified in the input prompt.
For the prompt reasoning method (e.g., T2I-R1 Jiang et al. (2025a)), as shown in Figure 1 (b), CoT is employed to analyze the original prompt from multiple perspectives, e.g., the subject, scene requirements, and other relevant aspects, in order to derive a more refined prompt. Prompt reasoning fully leverages the understanding capabilities of the unified model to enhance the initial prompt, and the resulting refined prompt is then used as the final input for image generation. This approach emphasizes the “understand first, generate later” pipeline, in which the understanding phase precedes the generation process. However, this pipeline focuses exclusively on language during the reasoning and lacks interaction with the generated images, which leads to a failure to capture the inherent limitations of the generation model. As illustrated in Figure 1 (b), although a refined prompt is obtained through reasoning, the process does not engage with generation, thereby preventing the recognition of generative weaknesses, e.g., the spatial relationship between the cup and the woman. As a result, the refined prompt fails to guide the reasoning to the correct direction.
@@ 2.1 Understanding-Generation Unified Model
Recent multimodal large language models Driess et al. (2023); Peng et al. (2023); Alayrac et al. (2022); Zhu et al. (2023b); Wang et al. (2024a); Bai et al. (2025); Chen et al. (2024); Li et al. (2023); Lin et al. (2023); Gao et al. (2023); Zhang et al. (2023) have demonstrated remarkable progress in visual understanding powered by the strong vision encoders Dosovitskiy et al. (2020); Radford et al. (2021); Girdhar et al. (2023); Zhu et al. (2023a); Guo et al. (2023); Lyu et al. (2024). Building on this success, unifying visual understanding and generation within a single framework has emerged as a central research frontier. Early unified models, such as Transfusion Zhou et al. (2024), Chameleon Team (2024), Emu3 Wang et al. (2024b), and Show-o Xie et al. (2024), employ visual tokenizers (e.g., VQ-VAE Van Den Oord et al. (2017)) to encode images into sequences of tokens, to enable seamless multimodal understanding and generation further. However, these discrete visual tokens introduce visual information loss, limiting the extraction of fine-grained semantic content. To address this, the Janus Chen et al. (2025b) series decouples visual encoding for understanding and generation by adopting separate encoders, though task conflicts within the shared LLM parameter space can hinder its performance. Meanwhile, BAGEL Deng et al. (2025) adopts a Mixture-of-Experts architecture, assigning autoregressive text generation and diffusion-based image synthesis to distinct components. Despite their effectiveness, these models still struggle to fully exploit their strong understanding capabilities during the generative process, which restricts their ability to produce images with complex logical content.
@@ 2.2 Chain-of-Thought for Text-to-Image Generation
Chain-of-Thought Wei et al. (2022) has played an effective role in LLM Xia et al. (2024); Zhang et al. (2024); Deng et al. (2024); Chen et al. (2025a); Kang et al. (2025); Wu et al. (2025b); Wang & Zhou (2024) and MLLM Wang et al. (2025); Xu et al. (2024); Shao et al. (2024); Zhao et al. (2025); Ma et al. (2025); Jiang et al. (2025b), enabling them to decompose complex tasks into structured intermediate steps. Recent research has extended CoT into text-to-image generation, primarily through verification-based and prompt-based approaches. Verification-based methods (e.g., ImageCoT Zhang et al. (2025a)) generate multiple candidate images via repeated sampling and then apply CoT to verify intermediate results. However, in this setting, reasoning functions solely as a verification stage, leaving the generative process unguided. Consequently, the generative capacity remains restricted to outcomes accessible through repeated sampling. In contrast, prompt-based methods (e.g., T2I-R1 Jiang et al. (2025a), ReasonGen-R1 Zhang et al. (2025b), and ImageGenCoT Liao et al. (2025)) leverage CoT to refine the input prompt by decomposing it into semantic aspects, to further improve the prompt for generation. However, this reasoning operates independently of the generation process and cannot overcome the limitations of generative capability.
@@ 3.2 The Understanding-in-Generation Reasoning
Figure 3: The comparison between the (a) Understanding-Generation-Separation and (b) Understanding-in-Generation reasoning.
@@ 3.4 Implementation
Base Model: We utilize BAGEL Deng et al. (2025) as the base unified model to demonstrate the effectiveness of our proposed UiG framework. BAGEL is a robust and versatile unified model designed with a Mixture-of-Experts architecture. It dedicates specialized components to autoregressive text generation and diffusion-based image generation. Notably, BAGEL has demonstrated superior performance among open-source unified models. To illustrate the significant improvement enabled by our UiG framework, we apply our method to this state-of-the-art open-source model.
@@ 4.1 Benchmark and Experiment Setup
Benchmark: We evaluate the performance of text-to-image reasoning using the TIIF benchmark Wei et al. (2025) and the WISE benchmark Niu et al. (2025). The TIIF benchmark provides a comprehensive framework for fine-grained assessment of text-to-image models. It features 36 novel prompt combinations spanning six compositional dimensions, along with 100 real-world, designer-level prompts that demand sophisticated aesthetic judgment. The WISE benchmark, a widely adopted standard in this domain, poses challenges by embedding world knowledge within prompts, thereby testing a model’s capacity for knowledge-based text-to-image reasoning.
Baselines: Our baseline models consist of original AR-based models, including Llamagen Sun et al. (2024), LightGen Wu et al. (2025a), Show-o Xie et al. , Infinity Han et al. (2025), Janus-Pro Chen et al. (2025b), Orthus Kou et al. (2024), Vila-u Wu et al. (2024), and Emu3 Wang et al. (2024b); as well as text-to-image reasoning models, including ImageCoT Zhang et al. (2025a), ReasonGen-R1 Zhang et al. (2025b), and T2I-R1 Jiang et al. (2025a).
@@ 4.3 Qualitative Results
As illustrated in Figure 4, we present a visual comparison between existing text-to-image reasoning methods (e.g., ImageCoT Zhang et al. (2025a), ReasonGen-R1 Zhang et al. (2025b), and T2I-R1 Jiang et al. (2025a)) and our proposed UiG framework. The visual results demonstrate a notable improvement in prompt alignment achieved by UiG in the generated images. For instance, given the prompt [”On a stool are four black balloons and three yellow bananas”], only UiG accurately generates the correct quantities for both balloons and bananas, as shown in Figure 4. Furthermore, in response to the prompt [”A wooden laptop is sitting silently on a wooden desk, not displaying a beautiful scenery”], all existing reasoning methods fail to generate an image without displaying scenery. In contrast, UiG produces a visually accurate result consistent with the prompt, as also shown in Figure 4. These visual comparisons show the substantial gains in prompt following and generation accuracy provided by our UiG framework. Additional results are presented in Section C.1 of the Appendix to further validate the effectiveness of UiG.
@@ 6 Conclusion
Limitation and Future Work: Our UiG is not available for video generative models. Therefore, future work will focus on improving the reasoning for video generation.
@@ B.1 More Details of Benchmarks
TIIF Benchmark is a comprehensive and fine-grained evaluation benchmark specifically designed to assess the instruction-following capabilities of modern Text-to-Image (T2I) models. Unlike prior benchmarks such as COMPBENCH++ Huang et al. (2025) or GENAI BENCH Li et al. (2024), which suffer from semantic redundancy, fixed prompt lengths, and coarse evaluation metrics, TIIF-Bench offers a diverse and hierarchically structured set of 5000 prompts that span a broad range of compositional and semantic complexity.
@@ D.2 Trade-off
As shown in Table 5, we analyze the trade-off between performance and latency across Verification-based reasoning, prompt reasoning, and our proposed UiG. To ensure a fair comparison, the analysis is conducted using the same base model (BAGEL Deng et al. (2025)). Our UiG achieves the best performance with lower latency in inference time compared to Verification-based reasoning, although it lags slightly behind prompt reasoning in terms of performance. These results demonstrate the overall advantage of UiG in balancing performance and latency.

### v3

@@ 1 Introduction
First, for verification-based reasoning (e.g., ImageCoT (Zhang et al., 2025a)), as illustrated in Figure 1 (a), this method constructs multiple generative branches through repeated sampling and assesses the generative potential throughout the generation process. Specifically, the state of each branch is evaluated to determine whether it still has the potential to continue generating. If the assessment gives a negative result, the branch is terminated. Here, the understanding of the unified model is applied exclusively to the verification of intermediate images, selecting the best output from among numerous samples. However, in this framework, the understanding capability is employed merely as a tool for validation and filtering, rather than providing effective guidance during the generation. As a result, the generative ability remains confined to scopes that can be reached through repeated sampling. As shown in Figure 1 (a), given the input prompt, [”There is a cup positioned behind the woman”], all outputs consistently place the cup in front of the woman. Consequently, even the best result obtained through verification-based reasoning fails to satisfy the spatial relationship specified in the input prompt.
For the prompt reasoning method (e.g., T2I-R1 (Jiang et al., 2025a)), as shown in Figure 1 (b), CoT is employed to analyze the original prompt from multiple perspectives, e.g., the subject, scene requirements, and other relevant aspects, in order to derive a more refined prompt. Prompt reasoning fully leverages the understanding capabilities of the unified model to enhance the initial prompt, and the resulting refined prompt is then used as the final input for image generation. This approach emphasizes the “understand first, generate later” pipeline, in which the understanding phase precedes the generation process. However, this pipeline focuses exclusively on language during the reasoning and lacks interaction with the generated images, which leads to a failure to capture the inherent limitations of the generation model. As illustrated in Figure 1 (b), although a refined prompt is obtained through reasoning, the process does not engage with generation, thereby preventing the recognition of generative weaknesses, e.g., the spatial relationship between the cup and the woman. As a result, the refined prompt fails to guide the reasoning to the correct direction.
@@ 2.1 Understanding-Generation Unified Model
Recent multimodal large language models (Driess et al., 2023; Peng et al., 2023; Alayrac et al., 2022; Zhu et al., 2023b; Wang et al., 2024a; Bai et al., 2025; Chen et al., 2024; Li et al., 2023; Lin et al., 2023; Gao et al., 2023; Zhang et al., 2023) have demonstrated remarkable progress in visual understanding powered by the strong vision encoders (Dosovitskiy et al., 2020; Radford et al., 2021; Girdhar et al., 2023; Zhu et al., 2023a; Guo et al., 2023; Lyu et al., 2024). Building on this success, unifying visual understanding and generation within a single framework has emerged as a central research frontier. Early unified models, such as Transfusion (Zhou et al., 2024), Chameleon (Team, 2024), Emu3 (Wang et al., 2024b), and Show-o (Xie et al., 2024), employ visual tokenizers (e.g., VQ-VAE (Van Den Oord et al., 2017)) to encode images into sequences of tokens, to enable seamless multimodal understanding and generation further. However, these discrete visual tokens introduce visual information loss, limiting the extraction of fine-grained semantic content. To address this, the Janus (Chen et al., 2025b) series decouples visual encoding for understanding and generation by adopting separate encoders, though task conflicts within the shared LLM parameter space can hinder its performance. Meanwhile, BAGEL (Deng et al., 2025) adopts a Mixture-of-Experts architecture, assigning autoregressive text generation and diffusion-based image synthesis to distinct components. Despite their effectiveness, these models still struggle to fully exploit their strong understanding capabilities during the generative process, which restricts their ability to produce images with complex logical content.
@@ 2.2 Chain-of-Thought for Text-to-Image Generation
Chain-of-Thought (Wei et al., 2022) has played an effective role in LLM (Xia et al., 2024; Zhang et al., 2024; Deng et al., 2024; Chen et al., 2025a; Kang et al., 2025; Wu et al., 2025b; Wang & Zhou, 2024) and MLLM (Wang et al., 2025; Xu et al., 2024; Shao et al., 2024; Zhao et al., 2025; Ma et al., 2025; Jiang et al., 2025b), enabling them to decompose complex tasks into structured intermediate steps. Recent research has extended CoT into text-to-image generation, primarily through verification-based and prompt-based approaches. Verification-based methods (e.g., ImageCoT (Zhang et al., 2025a)) generate multiple candidate images via repeated sampling and then apply CoT to verify intermediate results. However, in this setting, reasoning functions solely as a verification stage, leaving the generative process unguided. Consequently, the generative capacity remains restricted to outcomes accessible through repeated sampling. In contrast, prompt-based methods (e.g., T2I-R1 (Jiang et al., 2025a), ReasonGen-R1 (Zhang et al., 2025b), and ImageGenCoT (Liao et al., 2025)) leverage CoT to refine the input prompt by decomposing it into semantic aspects, to further improve the prompt for generation. However, this reasoning operates independently of the generation process and cannot overcome the limitations of generative capability.
@@ 3.2 The Understanding-in-Generation Reasoning
Figure 3: The comparison of reasoning processes between the (a) Understanding-Generation-Separation and (b) Understanding-in-Generation reasoning.
@@ 3.4 Implementation
Base Model: We utilize BAGEL (Deng et al., 2025) as the base unified model to demonstrate the effectiveness of our proposed UiG framework. BAGEL is a robust and versatile unified model designed with a Mixture-of-Experts architecture. It dedicates specialized components to autoregressive text generation and diffusion-based image generation. Notably, BAGEL has demonstrated superior performance among open-source unified models. To illustrate the significant improvement enabled by our UiG framework, we apply our method to this state-of-the-art open-source model.
@@ 4.1 Benchmark and Experiment Setup
Benchmark: We evaluate the performance of text-to-image reasoning using the TIIF benchmark (Wei et al., 2025) and the WISE benchmark (Niu et al., 2025). The TIIF benchmark provides a comprehensive framework for fine-grained assessment of text-to-image models. It features 36 novel prompt combinations spanning six compositional dimensions, along with 100 real-world, designer-level prompts that demand sophisticated aesthetic judgment. The WISE benchmark, a widely adopted standard in this domain, poses challenges by embedding world knowledge within prompts, thereby testing a model’s capacity for knowledge-based text-to-image reasoning.
Baselines: Our baseline models consist of original AR-based models, including Llamagen (Sun et al., 2024), LightGen (Wu et al., 2025a), Show-o (Xie et al., ), Infinity (Han et al., 2025), Janus-Pro (Chen et al., 2025b), Orthus (Kou et al., 2024), Vila-u (Wu et al., 2024), and Emu3 (Wang et al., 2024b); as well as text-to-image reasoning models, including ImageCoT (Zhang et al., 2025a), ReasonGen-R1 (Zhang et al., 2025b), and T2I-R1 (Jiang et al., 2025a).
@@ 4.3 Qualitative Results
As illustrated in Figure 4, we present a visual comparison between existing text-to-image reasoning methods (e.g., ImageCoT (Zhang et al., 2025a), ReasonGen-R1 (Zhang et al., 2025b), and T2I-R1 (Jiang et al., 2025a)) and our proposed UiG framework. The visual results demonstrate a notable improvement in prompt alignment achieved by UiG in the generated images. For instance, given the prompt [”On a stool are four black balloons and three yellow bananas”], only UiG accurately generates the correct quantities for both balloons and bananas, as shown in Figure 4. Furthermore, in response to the prompt [”A wooden laptop is sitting silently on a wooden desk, not displaying a beautiful scenery”], all existing reasoning methods fail to generate an image without displaying scenery. In contrast, UiG produces a visually accurate result consistent with the prompt, as also shown in Figure 4. These visual comparisons show the substantial gains in prompt following and generation accuracy provided by our UiG framework. Additional results are presented in Section C.1 of the Appendix to further validate the effectiveness of UiG.
@@ 6 Conclusion
Limitation and Future Work: Currently, our UiG does not support video generative models. Future work will focus on enhancing the reasoning capabilities for video generation.
@@ B.1 More Details of Benchmarks
TIIF Benchmark is a comprehensive and fine-grained evaluation benchmark specifically designed to assess the instruction-following capabilities of modern Text-to-Image (T2I) models. Unlike prior benchmarks such as COMPBENCH++ (Huang et al., 2025) or GENAI BENCH (Li et al., 2024), which suffer from semantic redundancy, fixed prompt lengths, and coarse evaluation metrics, TIIF-Bench offers a diverse and hierarchically structured set of 5000 prompts that span a broad range of compositional and semantic complexity.
@@ D.2 Trade-off
As shown in Table 5, we analyze the trade-off between performance and latency across Verification-based reasoning, prompt reasoning, and our proposed UiG. To ensure a fair comparison, the analysis is conducted using the same base model (BAGEL (Deng et al., 2025)). Our UiG achieves the best performance with lower latency in inference time compared to Verification-based reasoning, although it lags slightly behind prompt reasoning in terms of performance. These results demonstrate the overall advantage of UiG in balancing performance and latency.